OpenStudy (anonymous):

HELP PLEASEE ! ILL MEDAL ! The No-Crash Rental Company says the total cost to rent a dump truck can be found using the equation y = 45x + 75, where x represents the number of days that the truck is rented. How much would it cost to rent a truck for 5 days?

OpenStudy (anonymous):

Break the original equation down y = 45x + 75 45 * x = 225 x = number of days so 45 * 5 = 225 225 + 75 = 300 To rent the truck for 5 days it would cost 300$